OpenStudy (love_to_love_you):

Which option describes the left and right continuity of f(x) at x = 2? f(x) = {(x^3 + 1)^1/2 if 0 < x <= 2 { x + 1 if x > 2

OpenStudy (love_to_love_you):

only continuous from the right only continuous from the left continuous from both left and right continuous from neither

OpenStudy (sunnnystrong):

If:\[f(x)=\left\{\sqrt{ x^3+1} \right\} \] for x in the interval (0.2] & \[f(x)=\left\{x+1 \right\} \] for x in the interval (2, infinity) Than you can evaluate what happens to f(x) to the left & right of 2

OpenStudy (sunnnystrong):

is each equation continuous to the left & right of 2? @love_to_love_you

OpenStudy (love_to_love_you):

I think so